Assume we have the following PDF for the random variable X

x 0 1 2 3 4
f ( x ) .2 .1 .15 .3 .25

If Var ( X ) = 2.11, what is Var(3-2X)
Answer #1

Var(3-2X) = Var(3) + Var(2X)

                  = 0 + 22*Var(x)------------------------(Var(const) = 0 and Var(ax) = a2Var(X))

                  =4*2.11

                  = 8.44